OpenStudy (anonymous):

which are the solutions for the following equation. Check all that apply. (5x-2)^2=10

OpenStudy (anonymous):

A.\[A.x=\frac{ \sqrt10+2 }{ 5 }\]

OpenStudy (anonymous):

B.\[x=\frac{ \sqrt{12} }{ 5 }\]

OpenStudy (anonymous):

C.\[x=\sqrt{2}+2\]

OpenStudy (anonymous):

D.\[x=-\sqrt{2}+2\]

OpenStudy (anonymous):

E.\[x=\frac{ -\sqrt{12} }{ 5 }\]

OpenStudy (anonymous):

F.\[x=\frac{ -\sqrt{10}+2 }{ 5 }\]

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

uh yeah young money gettin rich girl u know wat is rack city b1tch rack rack city b1tch....

OpenStudy (anonymous):

\[(5x-2)^2=10 \]\[5x-2=\sqrt{10}\]\[5x=\sqrt{10}+2\]Then divide

OpenStudy (anonymous):

And why are you quoting lyrics?

OpenStudy (anonymous):

my mad i thought i was in the music section

OpenStudy (anonymous):

its A and F

OpenStudy (anonymous):

yup